# Classical simulation of non-Gaussian bosonic circuits

### v1

@@ 1 Introduction
If one is to believe in Feynman’s vision [1], understanding quantum dynamics may one day be a prime field of application of a working universal quantum computer. Without access to such a powerful computational tool, however, one has to resort to efficient, i.e., polynomial time, classical methods that can be used to study the behavior of certain quantum many-body systems evolving e.g., under unitaries and measurement. A difficulty here is to identify classes of quantum computations which allow for efficient classical simulation algorithms. Well-known examples include matchgate circuits [2, 3], Gaussian dynamics of either fermions [4, 5] or bosons, and Clifford circuits (stabilizer computations) [6]. For each of these classes of quantum dynamics, polynomial-time simulation methods have been constructed.
@@
For bosonic systems consisting of say $n$ modes, low-order moments of the mode operators of a state (that is, expectation values of constant-degree polynomials of the quadratures) can easily be kept track of under Gaussian dynamics (in particular, Gaussian unitaries and measurement): This requires only a polynomial-time computation, and implies that (bosonic) linear optics, i.e., computations with such operations starting from an initial Gaussian state, can be simulated efficiently. This is the case for both strong simulation (where the goal is to compute the output probability for a given measurement outcome) as well as weak simulation (where the goal is to sample from the output distribution defined by measurement), assuming the classical algorithm has access to independent unit-variance centered Gaussian random variables.
@@
While the framework of quantum linear optics is amenable to efficient classical simulation algorithms, it is not sufficient to describe a number of information-processing applications of interest. Consider for example quantum error correction. To quote from the abstract of a pioneering paper [7], “[…] Gaussian operations are of no use for protecting Gaussian states against Gaussian errors […]” (see [8] for another no-go result). It is thus necessary to augment quantum optics by non-Gaussian resource states. For instance, displacement errors can be protected against by encoding a qubit into an (approximate) Gottesman-Kitaev-Preskill (GKP) code [9]. Other proposals for protecting against displacement errors include the family of oscillator-to-oscillator codes [10], where encoding proceeds by applying a Gaussian unitary to a bosonic mode, with a collection of auxiliary systems prepared in GKP-states. To protect against photon loss (another ubiquitous example of Gaussian errors), coherent cat codes [11, 12, 13, 14] have been developed.
@@
What do GKP-, oscillator-to-oscillator- and coherent cat-codes have in common? Most importantly, they are experimentally feasible, and corresponding schemes have been realized in the lab. What makes this possible is the simple structure of the associated circuits: While the initial states are non-Gaussian, subsequent operations (e.g., for applying logical gates) typically only involve Gaussian unitaries. Second, analyzing their properties presents a difficult challenge for theory. In particular, unlike for Gaussian initial state, it is unclear in general how to construct classical simulation algorithms which work for non-Gaussian initial states. It is clear that the hardness, i.e., complexity of classical simulation is a function of these states. Our goal is to quantify this, i.e., we ask: What is the computational effort required to (strongly) simulate a Gaussian quantum circuit on a non-Gaussian input state?
@@
In an effort to go beyond this framework of linear optics, several alternative methods for simulating bosonic dynamics have been developed. For example, it has been shown that if a circuit of Gaussian unitaries is applied to an initial state with a positive Wigner function, then it is possible to efficiently sample from the output distribution of a subsequent measurement [15, 16]. Since Gaussian states have positive Wigner function, this constitutes a strict generalization of the covariance matrix formalism to a larger class of states. One way of putting this result is to say that Wigner function negativity is a precondition for a quantum advantage [17].
@@
In more recent work [18], a different set of non-Gaussian initial states for subsequent Gaussian operations have been shown to also lead to a class of computations that can be simulated efficiently in the strong sense. The authors of Ref. [18] give an algorithm which works for initial pure states whose coefficients in the Fock basis have bounded support. Their simulation algorithms have polynomial runtime in the size of this support, as well as exponential runtime in the mean photon number of the initial state.
@@
For GKP- and cat-codes (and more generally when dealing with infinite-dimensional Hilbert spaces), perhaps the most widely used approach to numerical simulation (see e.g., [19]) is to project onto a finite-dimensional Hilbert space (and subsequently do, e.g., state vector simulation). Concretely, e.g., in the work [19], this is done by replacing GKP- and cat-states by finite Fock state superpositions $\sum_{n=0}^{N_{\max}}c_{n}|n\rangle$ (with $N_{\max}$ as large as manageable). While this is sufficient for the analysis of single-mode bosonic codes, it is desirable to find better approaches that may be more scalable. In this context, the consideration of number (Fock) states also does not appear to be ideal as these have non-trivial dynamics under Gaussian operations. In fact, GKP- and cat-states have an important commonality that should be exploited: These are all states that are superpositions of (well-understood) Gaussian states. Motivated by this, we refine our goal and ask: can the set of initial states to a Gaussian quantum circuit be extended in a different direction beyond Gaussian states, namely superpositions of Gaussian states, while maintaining efficient (strong) simulability?
@@
In more detail, consider the following type of $n$-mode bosonic quantum computation, which augments linear optics by the use of an initial state $\Psi$ form a certain set of states $\mathcal{S}\subset L^{2}(\mathbb{R})^{\otimes n}$. Here and throughout we use the notation []=$[\ell]=\{1,\ldots,\ell\}$ for $\ell\in\mathbb{N}$.
@@
    For each $t\in\{1,\ldots,s\}$, apply a Gaussian unitary $U^{(t)}$ from the set of all Gaussian unitaries that are either
@@
where we write $|\alpha\rangle=|\alpha_{1}\rangle\otimes\cdots\otimes|\alpha_{k}\rangle$ for a product of coherent states. Since the set of unitaries considered in Step 2 generates all Gaussian unitaries, this class of computations encompasses the three mentioned (known) examples – the set of Gaussian computations (with initial state $\Psi$ belonging to the set $\mathcal{S}=\mathrm{Gauss}_{n}$ of pure Gaussian states), as well as the extension by initial states with positive Wigner functions, and the set of computations starting from states with finite support in the Fock-basis and bounded photon number (with $\mathcal{S}$ defined correspondingly) – as well as the direction taken here, i.e., computations starting from superpositions of Gaussian states.
2 Our contribution
@@
Here the runtime of the algorithm $\mathsf{SimulateExactly}$ is defined as the number of elementary arithmetic operations over the reals, and we assume that we have arbitrary-precision arithmetic available. This precision is needed since we are not making any assumptions about the initial state $\Psi$ (except for the fact that it is a linear combination of Gaussian states with $\chi$ terms): The algorithm  $\mathsf{SimulateExactly}$ involves real numbers that are essentially the inverse of the energy (mean photon number) of $\Psi$. In particular, if no energy bound is imposed, these scalars can become arbitrarily small.
@@
$\mathcal{S}^{\mathrm{Gauss}_{n}}_{N}(\chi)$ :=$:=\left\{\Psi\in\mathcal{S}^{\mathrm{Gauss}_{n}}(\chi)\ \big{|}\ \langle\Psi,\sum_{j=1}^{\chi}a_{j}^{\dagger}a_{j}\Psi\rangle\leq N\right\}$ (3)
@@
The computational complexity of our algorithms is tied to the (bosonic) Gaussian rank of the initial state $\Psi$, i.e., the minimal integer $\chi\in\mathbb{N}$ such that $\Psi\in\mathcal{S}^{\mathrm{Gauss}_{n}}(\chi)$. This definition is analogous to the stabilizer rank [20, 20], as well as to the fermionic Gaussian rank [21, 22]. Assuming that the initial state is provided as a superposition of Gaussian states with the minimum number $\chi$ of terms, the algorithms $\mathsf{SimulateExactly}$ and $\mathsf{SimulateApproximately}$ scale quadratically and linearly in $\chi$, respectively. Starting from this observation, one may introduce the notion of the bosonic Gaussian extent of $\Psi$, similar to the way the stabilizer [23] and fermionic Gaussian extent [21, 22, 24] are defined. This quantity is typically much smaller than the Gaussian rank and is a proxy for an approximate notion of Gaussian rank. The runtime of an appropriately modified version of the algorithm $\mathsf{SimulateApproximately}$ scales linearly in this quantity.
@@
Classical strong simulation is indeed stronger than weak simulation in the sense that an algorithm for strong simulation gives an algorithm for weak simulation, provided that there is an efficient algorithm to sample from the output probability distribution associated to measurement [25]. This statement carries over to continuous variable computations by discretizing the output probability density, provided there is an efficient binning of the space of outcomes and the discretized probabilities can be computed efficiently from the output probability density and sampled from efficiently [18]. Then, if provided, an efficient binning of the outcome space as well as efficient algorithms to compute and sample discretized probabilities from the output probability density augment our classical simulation methods in a way that allows weak simulation.
@@
Our work mirrors earlier findings in the context of stabilizer [20, 23, 26], matchgate and fermionic linear optics computations [21, 24]. It makes an experimentally important new class of non-Gaussian circuits accessible to numerical simulation.
@@ 2.1 Technical building blocks
where $\alpha\in\mathbb{C}^{n}$ specifies a displacement operator $D(\alpha)$, and where $\rho(\Gamma,d)$ denotes a Gaussian state with covariance matrix $\Gamma$ and displacement vector $d$. Using this formula, which is a generalization of an expression previously derived in [27] for the case of centered Gaussian states and $\alpha$ set to zero, one can compute expressions of the form
@@
for three Gaussian states $\psi_{j}$, $j=1,2,3$, in terms of their covariance matrices and displacement vectors, see Lemma 4.4. Our construction then relies on using a description of Gaussian states (including their global phase) that can be combined with this formula to evaluate inner products (including phases). We have previously used such an approach in the context of non-Gaussian operations using fermions [21].
@@ 3 Background
$[a_{j},a_{k}^{\dagger}]=\delta_{j,k}I\qquad,\qquad[a_{j},a_{k}]=[a_{j}^{\dagger},a_{k}^{\dagger}]=0\qquad\text{for}\qquad j,k\in[n]\ .$ (8)
A number state is defined as
$|k\rangle=(a_{1}^{\dagger})^{k_{1}}\cdots(a_{n}^{\dagger})^{k_{n}}|0\rangle\qquad\text{for}\qquad k\in\mathbb{N}_{0}=\mathbb{N}\cup\{0\}\ ,$ (9)
@@
$a_{j}|0\rangle=0\qquad\text{for all}\qquad j\in[n]\ .$ (10)
@@
$Q_{j}&=(a_{j}+a_{j}^{\dagger})/\sqrt{2}\\ P_{j}&=-i(a_{j}-a_{j}^{\dagger})/\sqrt{2}\end{aligned}\end{gathered}\qquad\text{for}\qquad j\in[n]$ (12)
@@
$\Omega=\bigoplus_{j=1}^{n}\begin{pmatrix}0&1\\ -1&0\end{pmatrix}$ (15)

### v2

@@ 1 Introduction
If one is to believe in Feynman’s vision [feynmanSimulatingPhysicsComputers1982], understanding quantum dynamics may one day be a prime field of application of a working universal quantum computer. Without access to such a powerful computational tool, however, one has to resort to efficient, i.e., polynomial time, classical methods that can be used to study the behavior of certain quantum many-body systems evolving e.g., under unitaries and measurement. A difficulty here is to identify classes of quantum computations which allow for efficient classical simulation algorithms. Well-known examples include matchgate circuits [doi:10.1137/S0097539700377025, Jozsa_2008], Gaussian dynamics of either fermions [PhysRevA.65.032325, knill2001fermionic] or bosons [WANG_2007], and Clifford circuits (stabilizer computations) [PhysRevA.70.052328]. For each of these classes of quantum dynamics, polynomial-time simulation methods have been constructed.
@@
For bosonic systems consisting of say $n$ modes, low-order moments of the mode operators of a state (that is, expectation values of constant-degree polynomials of the quadratures) can easily be kept track of under Gaussian dynamics (in particular, Gaussian unitaries and measurement): This requires only a polynomial-time computation, and implies that (bosonic) linear optics, i.e., computations with such operations starting from an initial Gaussian state can be simulated efficiently. This is the case for both strong simulation (where the goal is to compute the output probability for a given measurement outcome) as well as weak simulation (where the goal is to sample from the output distribution defined by measurement), assuming the classical algorithm has access to independent unit-variance centered Gaussian random variables.
@@
While the framework of quantum linear optics is amenable to efficient classical simulation algorithms, it is not sufficient to describe a number of information-processing applications of interest. Consider for example quantum error correction. To quote from the abstract of a pioneering paper [NisetFiurasekCerf09], “[…] Gaussian operations are of no use for protecting Gaussian states against Gaussian errors […]” (see [GKPToric] for another no-go result). It is thus necessary to augment quantum optics by non-Gaussian resource states. For instance, displacement errors can be protected against by encoding a qubit into an (approximate) Gottesman-Kitaev-Preskill (GKP) code [gkp]. Other proposals for protecting against displacement errors include the family of oscillator-to-oscillator codes [nohoscillatortooscillator], where encoding proceeds by applying a Gaussian unitary to a bosonic mode, with a collection of auxiliary systems prepared in GKP-states. To protect against photon loss (another ubiquitous example of Gaussian errors), coherent cat codes [PhysRevA.94.042332, PhysRevLett.119.030502, PhysRevLett.111.120501, allopticalcatcode] have been developed.
@@
What do GKP-, oscillator-to-oscillator- and coherent cat-codes have in common? Most importantly, they are experimentally feasible, and corresponding schemes have been realized in the lab. What makes this possible is the simple structure of the associated circuits: While the initial states are non-Gaussian, subsequent operations (e.g., for applying logical gates) typically only involve Gaussian unitaries. Second, analyzing their properties presents a difficult challenge for theory. In particular, unlike for Gaussian initial states, it is unclear in general how to construct classical simulation algorithms which work for non-Gaussian initial states. It is clear that the hardness, i.e., complexity of classical simulation is a function of these states. Our goal is to quantify this, i.e., we ask: What is the computational effort required to (strongly) simulate a Gaussian quantum circuit on a non-Gaussian input state?
1.1 Prior work
@@
In an effort to go beyond this framework of linear optics, several alternative methods for simulating bosonic dynamics have been developed. For example, it has been shown that if a circuit of Gaussian unitaries is applied to an initial state with a positive Wigner function, then it is possible to efficiently sample from the output distribution of a subsequent measurement [Veitch_2013, PhysRevLett.109.230503]. Since Gaussian states have a positive Wigner function, this constitutes a strict generalization of the covariance matrix formalism to a larger class of states. One way of putting this result is to say that Wigner function negativity is a precondition for a quantum advantage [Veitch_2012]. Along similar lines, Ref. [PRXQuantum.2.040315] considered non-Gaussian states whose Wigner function can be written as a linear combination of Gaussian Wigner functions. Several natural non-Gaussian states were shown to be describable in this way. When using complex Gaussians, this framework can also describe superpositions of pure Gaussian states. Update rules for such descriptions under Gaussian operations were established.
@@
In more recent work [PhysRevResearch.3.033018, PhysRevLett.130.090602], a different set of non-Gaussian initial states for subsequent Gaussian operations have been shown to also lead to a class of computations that can be simulated efficiently in the strong sense. The authors of Refs. [PhysRevResearch.3.033018, PhysRevLett.130.090602] give algorithms which work for initial pure states whose coefficients in the Fock basis have bounded support. Their simulation algorithms have polynomial runtime in the size of this support. Additionally, the algorithms proposed in Refs. [PhysRevResearch.3.033018] and [PhysRevLett.130.090602] have exponential runtime respectively in the mean photon number of the initial state and in a non-Gaussianity measure called the stellar rank.
Another instance of an efficient classically simulable computation put forward in Refs. [Calcluth2022efficientsimulation, PhysRevA.107.062414] includes infinitely squeezed GKP-states together with a certain set of Gaussian unitary operations and homodyne measurements. In subsequent work [PhysRevA.107.062414, calcluth2024sufficient], the authors establish that certain resourceful/magic states (such as the vacuum state) can promote such a computation to universality.
@@
For GKP- and cat-codes (and more generally when dealing with infinite-dimensional Hilbert spaces), perhaps the most widely used approach to numerical simulation (see e.g., [performancestructurebosonic]) is to project onto a finite-dimensional Hilbert space (and subsequently do, e.g., state vector simulation). Concretely, e.g., in the work [performancestructurebosonic], this is done by replacing GKP- and cat-states by finite Fock state superpositions $\sum_{n=0}^{N_{\max}}c_{n}|n\rangle$ (with $N_{\max}$ as large as manageable). While this is sufficient for the analysis of single-mode bosonic codes, it is desirable to find better approaches that may be more scalable. In this context, the consideration of number (Fock) states also does not appear to be ideal as these have non-trivial dynamics under Gaussian operations. An alternative approach proposed in Ref. [Marshall:23] considers Gaussian unitary evolution and Fock basis measurements applied to non-Gaussian states decomposed into a superposition of coherent states, defining a notion called coherent state rank. These algorithms are shown to be more efficient than algorithms considering Fock state superpositions.
1.2 Towards new simulation algorithms
Gaussian states and operations define a well-known class of many-body quantum dynamics that is amenable both to analytical considerations, as well as efficient classical simulation. Motivated by this, we ask: can the set of initial states to a Gaussian quantum circuit be extended in a different direction beyond Gaussian states, namely superpositions of Gaussian states, while maintaining efficient (strong) simulability?
@@ 1.2 Towards new simulation algorithms
In more detail, consider the following type of $n$-mode bosonic quantum computation, which augments linear optics by the use of an initial state $\Psi$ form a certain set of states $\mathcal{S}\subset L^{2}(\mathbb{R})^{\otimes n}$. Here and throughout we use the notation $\ell]=\{1,\ldots,\ell\}$ for $\ell\in\mathbb{N}$.
@@
    For each $t\in\{1,\ldots,s\}$, apply a unitary $U^{(t)}$ from one of the following Gaussian unitaries
@@
where we write $|\alpha\rangle=|\alpha_{1}\rangle\otimes\cdots\otimes|\alpha_{k}\rangle$ for a product of coherent states. Since the set of unitaries considered in Step 2 generates all Gaussian unitaries, this class of computations encompasses the mentioned (known) examples – the set of Gaussian computations (with initial state $\Psi$ belonging to the set $\mathcal{S}=\mathrm{Gauss}_{n}$ of pure Gaussian states), as well as the extension by initial states with positive Wigner functions, the set of computations starting from states with finite support in the Fock-basis and bounded photon number (with $\mathcal{S}$ defined correspondingly), and the set of computations starting from superpositions of coherent states – as well as the direction taken here, i.e., computations starting from superpositions of Gaussian states.
1.3 Our contribution
@@
The constant in the $O(\cdot)$-notation depends on the Gaussian unitaries $\{U^{(t)}\}_{t=1}^{s}$ applied, see Section 3.3. Here the runtime of the algorithm $\mathsf{SimulateExactly}$ is defined as the number of elementary arithmetic operations over the reals, and we assume that we have arbitrary-precision arithmetic available. This precision is needed since we are not making any assumptions about the initial state $\Psi$ (except for the fact that it is a linear combination of Gaussian states with $\chi$ terms): The algorithm $\mathsf{SimulateExactly}$ involves real numbers that are essentially the inverse of the energy (mean photon number) of $\Psi$. In particular, if no energy bound is imposed, these scalars can become arbitrarily small.
@@
$\mathcal{S}^{\mathrm{Gauss}_{n}}_{N}(\chi)$ $:=\left\{\Psi\in\mathcal{S}^{\mathrm{Gauss}_{n}}(\chi)\ \big|\ \langle\Psi,\sum_{j=1}^{\chi}a_{j}^{\dagger}a_{j}\Psi\rangle\leq N\right\}$ (4)
@@
The computational complexity of our algorithms is tied to the (bosonic) Gaussian rank of the initial state $\Psi$, i.e., the minimal integer $\chi\in\mathbb{N}$ such that $\Psi\in\mathcal{S}^{\mathrm{Gauss}_{n}}(\chi)$. This definition is analogous to the stabilizer rank [bravyiImprovedClassicalSimulation2016, bravyiImprovedClassicalSimulation2016], as well as to the fermionic Gaussian rank [dias2023classical, cudby2023gaussian]. Assuming that the initial state is provided as a superposition of Gaussian states with the minimum number $\chi$ of terms, the algorithms $\mathsf{SimulateExactly}$ and $\mathsf{SimulateApproximately}$ scale quadratically and linearly in $\chi$, respectively. Starting from this observation, one may introduce the notion of the bosonic Gaussian extent of $\Psi$, similar to the way the stabilizer [bravyiSimulationQuantumCircuits2019a] and fermionic Gaussian extent [dias2023classical, cudby2023gaussian, reardon-smithImprovedSimulationQuantum2023] are defined. This quantity is typically much smaller than the Gaussian rank and is a proxy for an approximate notion of Gaussian rank. The runtime of an appropriately modified version of the algorithm $\mathsf{SimulateApproximately}$ scales linearly in this quantity.
@@
Classical strong simulation is indeed stronger than weak simulation in the sense that an algorithm for strong simulation gives an algorithm for weak simulation, provided that there is an efficient algorithm to sample from the output probability distribution associated to measurement [Pashayan2020fromestimationof]. This statement carries over to continuous variable computations by discretizing the output probability density, provided there is an efficient binning of the space of outcomes and the discretized probabilities can be computed efficiently from the output probability density and sampled from efficiently [PhysRevResearch.3.033018]. Then, if provided, an efficient binning of the outcome space as well as efficient algorithms to compute and sample discretized probabilities from the output probability density augment our classical simulation methods in a way that allows weak simulation.
@@
Our work mirrors earlier findings in the context of stabilizer [bravyiImprovedClassicalSimulation2016, bravyiSimulationQuantumCircuits2019a, pashayanFastEstimationOutcome2022b], matchgate and fermionic linear optics computations [dias2023classical, reardon-smithImprovedSimulationQuantum2023]. It makes an experimentally important class of non-Gaussian circuits accessible to numerical simulation.
1.4 Application of our algorithm to bosonic error correction
We note that there is by now a wide variety of classical simulation algorithms for bosonic dynamics (see the discussion in Section 1.1). Since they typically apply to different sets of (initial) states and operations, a direct comparison of their merits is non-trivial. Rather, the algorithm best suited for classical simulation heavily depends on the application considered.
As already mentioned, we believe that an area of application particularly suited for our algorithm is that of bosonic error correction. Concretely, an (approximate) Gottesman-Kitaev-Preskill (GKP) code [gkp] encodes a qubit into the two-dimensional subspace of an oscillator defined by the two states
$|\overline{b}\rangle_{\mathrm{GKP}}$ $\propto\sum_{s=-\infty}^{\infty}e^{-\frac{\kappa^{2}}{2}\left((2s+1)\sqrt{2\pi}\right)^{2}}e^{-i(2s+b)P\sqrt{2\pi}}\left|\psi_{\Delta}\right\rangle\ ,$ (6)
for $b\in\{0,1\}$, where $\psi_{\Delta}(x)=e^{-x^{2}/(2\Delta^{2})}/(\pi\Delta^{2})^{1/4}$ is a Gaussian wavefunction, and where $(\Delta,\kappa)$ describes the shape (squeezing) of the GKP-state, see Fig. 1. This code protects against (small) random displacement errors in phase space. Fault-tolerant quantum computation with a CV system can be realized by means of copies of these GKP-states and Gaussian operations: Indeed, GKP-states and such operations provide computational universality [PhysRevLett.123.200502].
Because of the Gaussian envelope in Eq. (6) (see Fig. 1), GKP-states can be approximated well by a finite linear combination of Gaussian states. The behavior of a fault-tolerant quantum circuit with an initial state that is a tensor product of states of the form (6) can therefore be studied using our algorithm.
Similar considerations apply to so-called coherent cat-states, which have been proposed for protecting against photon loss (another ubiquitous example of Gaussian errors), see e.g., [PhysRevA.94.042332, PhysRevLett.119.030502, PhysRevLett.111.120501, allopticalcatcode]. Here a qubit is encoded into the span of the two states
$|\overline{b}\rangle_{\mathrm{CAT}}$ $\propto|i^{b}\alpha\rangle+|i^{b}(-\alpha)\rangle\ ,$ (7)
where $b\in\{0,1\}$. Since the states (7) only involve two Gaussian states in the superposition (i.e., have Gaussian rank $2$), they are even more compatible with our simulation method.
Another related example of a circuit simulable with our approach is the encoding circuit of an oscillator-to-oscillator code [nohoscillatortooscillator]: Here the encoding proceeds by applying a Gaussian unitary to a bosonic mode, with a collection of auxiliary systems prepared in GKP-states, see Fig. 2. Our circuit applies whenever the state to be encoded is itself decomposed into a finite linear combination of Gaussian states. We note that such corresponding states and circuit have been implemented experimentally, see e.g., [campagne-ibarcqQuantumErrorCorrection2020].
[…]
@@ 1.5 Technical building blocks
where $\alpha\in\mathbb{C}^{n}$ specifies a displacement operator $D(\alpha)$, and where $\rho(\Gamma,d)$ denotes a Gaussian state with covariance matrix $\Gamma$ and displacement vector $d$. Using this formula, which is a generalization of an expression previously derived in [PhysRevA.61.022306] for the case of centered Gaussian states and $\alpha$ set to zero, one can compute expressions of the form
@@
for three Gaussian states $\psi_{j}$, $j=1,2,3$, in terms of their covariance matrices and displacement vectors, see Lemma 3.4. Our construction then relies on using a description of Gaussian states (including their global phase) that can be combined with this formula to evaluate inner products (including phases). We have previously used such an approach in the context of non-Gaussian operations using fermions [dias2023classical]. The use of expressions of the form (9) allows us to keep track of the relative phases between Gaussian states in a superposition. We note that there may be other ways of achieving this. For example, Ref. [yao2023design] provides formulas for the global phase of a composition of Gaussian unitaries, which could serve this purpose. Such an approach has been pursued in the fermionic context by Ref. [reardon-smithImprovedSimulationQuantum2023].
@@ 2 Background
$[a_{j},a_{k}^{\dagger}]=\delta_{j,k}I\quad,\quad[a_{j},a_{k}]=[a_{j}^{\dagger},a_{k}^{\dagger}]=0$ (11)
for $j,k\in[n]$. A number state is defined as
$|k\rangle=(a_{1}^{\dagger})^{k_{1}}\cdots(a_{n}^{\dagger})^{k_{n}}|0\rangle\quad\text{for}\quad k\in\mathbb{N}_{0}=\mathbb{N}\cup\{0\}\ ,$ (12)
@@
$a_{j}|0\rangle=0\quad\text{for all}\quad j\in[n]\ .$ (13)
@@
$\begin{gathered}\begin{aligned} Q_{j}&=(a_{j}+a_{j}^{\dagger})/\sqrt{2}\\ P_{j}&=-i(a_{j}-a_{j}^{\dagger})/\sqrt{2}\end{aligned}\end{gathered}\quad\text{for}\quad j\in[n]$ (15)
@@
$\Omega=\bigoplus_{j=1}^{n}\begin{pmatrix}0&1\\ -1&0\end{pmatrix}\$(18)